OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (6, 2) and (4, 8)? y = ? + ?

OpenStudy (anonymous):

@Mehek14 @amorfide @iambatman

OpenStudy (anonymous):

@peterg2001 @Tayeni

OpenStudy (anonymous):

? ye Im here

OpenStudy (anonymous):

let me answer tht if i can

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

Hmm so wht exactly do u mean, is they asking like the difference that is turned to a adding equation?

OpenStudy (anonymous):

use y = mx+b the two points (6,2) and (4,8) give you the equatioins 2 = m*6+b 8 = m*4+b take the first equation minus the second to get -6 = 2m m= -3 plug m=-3 into either of the original equations to get 2 = -3*6+b b=20 so your linear equation will be y = -3*x + 20

OpenStudy (anonymous):

they want to know the equation

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

thx @dagrothus

OpenStudy (anonymous):

u to0 @Tayeni

OpenStudy (anonymous):

i was about to say maybe-3x + 12 = y but i see i was totally off